# Quantum Clocks Tick Faster: Entanglement, Contextuality, and the Flow of Time

### v1

@@ Abstract
Building on the recent proposal that a single “bona fide” clock suffices to define spacetime’s metric, we introduce an Entangled Clock protocol based on singlet-state correlations. Invoking Zeilinger’s Foundational Principle, we argue that while the local flow of time, operationally defined as a sequence of detector “ticks,” is irreducibly random (one bit per elementary system), the synchronized flow between spatially separated observers depends on their measurement geometry. Comparing the quantum prediction for the coincidence rate with Peres’ classical “bomb fragment” model, we find that at obtuse relative angles ($\theta\approx 140^{\circ}$) the entangled clock exhibits a ${\sim}13\%$ higher synchronized tick rate than this linear classical benchmark. This “temporal acceleration” is linked to contextuality: following Peres, “unperformed experiments have no results,” and quantum systems are not constrained to maintain consistency with all counterfactual measurement settings. We stress, however, that for any single measurement angle a suitably tailored classical model can reproduce the quantum rate. The genuinely nonclassical character of the entangled clock emerges only when correlations at several angles are considered simultaneously and are shown to violate Bell-type inequalities. In this sense, the violation of Bell-type bounds serves as a certification that the shared time standard is genuinely quantum.
@@ VI Experimental Considerations
VII Conclusion
We have proposed an operational protocol in which spatially separated quantum clocks, synchronized via singlet-state correlations, exhibit a measurable enhancement in their coincidence rate compared to Peres-type classical mechanisms based on conserved angular momentum and, more generally, to any local realistic model once correlations at several measurement angles are taken into account. By unifying the “single-unit universe” concept with an information-theoretic view of quantum randomness, we showed that the “bona fide clock” fixing the metric scale is locally random but relationally structured.
At a relative measurement angle of $\theta\approx 140^{\circ}$, this “temporal speedup” in the synchronized tick rate reaches ${\sim}13\%$ relative to the linear bomb-model benchmark. The effect traces to contextuality: quantum systems need not maintain consistency with unperformed measurements, freeing the correlation function from the linear constraints of specific hidden-variable models. Certification of this quantum advantage requires Bell-type inequality tests across multiple measurement contexts—similar to private randomness [14]—confirming that the cosine modulation cannot be replicated by any local realistic mechanism. In this sense, the entangled clock embodies the informational character of quantum time: random in isolation, relational in synchrony, and—when judged against the full set of classical constraints—faster than classical physics permits.

### v2

# Certified Private Quantum Time Ticks Away Faster than Any Classical Clock

@@ Abstract
We introduce the concept of an entangled clock, where the flow of time is operationally defined by the discrete registration of measurement outcomes on a singlet state. Comparing the synchronization rate of two such clocks against classical models, we find that at obtuse relative angles, the quantum clock ticks strictly faster than this classical benchmark. We further propose a protocol for Certified Private Time, adapting device-independent randomness certification to the temporal domain; this guarantees a sovereign timeline that ticks away faster than any local realistic mechanism allows.
@@ VI Experimental Considerations
VII Certified Private Time
The discussion above highlights a crucial distinction: while a classical model can be engineered to reproduce quantum statistics for a single context, it fails to reproduce the full correlation structure across multiple contexts. This brings us to a stronger application of the entangled clock: the certification of “private time.”
As reviewed by Acín and Masanes [14], the violation of Bell inequalities serves as a device-independent certificate for randomness. In the context of random number generation (QRNG), this certification ensures that the output bits were not pre-determined by any hidden variable $\lambda$ available to an adversary (Eve). Transposing this logic to metrology, we can conceive of a protocol for Certified Private Time.
VII.1 The Memory-Stick Clock
Consider the temporal equivalent of the “memory-stick attack” described in Ref. [14]. An adversary, Eve, manufactures two clocks and gives them to Alice and Bob. These clocks are classical devices containing a pre-recorded sequence of ticks derived from a hidden variable $\lambda$. To Alice and Bob, the ticks appear stochastic (e.g., mimicking radioactive decay), and they may even exhibit correlations. However, because the sequence is pre-determined, Eve possesses a “lookup table” for time. She knows exactly when Alice’s clock will tick before it happens. In such a scenario, Alice’s time is not her own; it is fully correlated with the adversary’s timeline.
In a classical universe (or one governed by local hidden variables), one cannot prove that a clock is not a playback device. The flow of time could be merely the reading of a script written at the Big Bang (super-determinism aside).
VII.2 Protocol for Time Certification
The entangled clock allows Alice and Bob to rule out this “playback” scenario. By switching between measurement settings (angles $\vec{a},\vec{a}^{\prime}$ for Alice and $\vec{b},\vec{b}^{\prime}$ for Bob) and accumulating the statistics of their coincident ticks, they can estimate the CHSH parameter $S$.
If they observe $S>2$, they certify that the joint state does not admit a local realistic description. Physically, this implies that the specific outcomes (the ticks) did not exist prior to the measurement events. Consequently, the time signal generated by the clock is:
  1. 1.
    Intrinsic: The ticks are generated in situ by the act of measurement, not read from a memory.
  2. 2.
    Private: Because the outcomes are fundamentally indeterministic, they are uncorrelated with any external system, including Eve’s variables.
VII.3 Escaping the Block Universe
This leads to a radical reinterpretation of the “flow” of time. In a standard operational view, time is a parameter inferred from correlations between a clock system and the rest of the universe. However, if the clock’s ticks are certified random via Bell violation, the clock is dynamically decoupled from the rest of the universe until the moment of the tick.
A “Certified Private Clock” thus ticks at a rate that is not only statistically distinct (the “faster” rate at obtuse angles discussed in Sec. IV) but ontologically distinct. It generates a local history that no external observer could have predicted. Just as device-independent protocols use Bell violation to expand a short private random seed into a distinct, private key, the entangled clock expands a short interval of private operational agency (the choice of settings) into a sustained, private flow of time.
This certification is robust against the “middleman attack” often cited in foundational debates. Even if Eve builds the source and controls the channels, she cannot force the devices to violate the CHSH inequality with pre-programmed ticks unless she knows the measurement settings (the context) in advance. To prevent this, Alice and Bob must ensure “measurement (parameter)independence” [15]. This can be achieved either through genuine free choice or, alternatively, by consuming a short, predefined private random key to govern the settings. In the latter case, the system functions as a temporal implementation of Acín’s expansion protocol [14], growing the initial finite seed into a sustained, sovereign timeline that is mathematically guaranteed to be uncorrelated with any master clock Eve might hold.
VIII Conclusion
We have proposed an operational protocol for Certified Private Time, in which spatially separated quantum clocks are synchronized via singlet-state correlations. By adapting device-independent certification methods [14], we showed that if observers consume a private seed to ensure measurement independence, the resulting timeline is certified to be “sovereign”—generated in situ and mathematically uncorrelated with any pre-existing external script.
Crucially, these quantum clocks differ from classical counterparts not just in their security, but in their rate. For instance, at obtuse relative angles ($\theta\approx 140^{\circ}$), the synchronized tick rate exhibits a “temporal speedup” of ${\sim}13\%$ relative to Peres’ classical bomb-fragment model. This enhancement obeys the conservation of probability: the quantum correlation function effectively “squeezes” the statistics, populating the synchronized states (++++ and --) more densely than the linear classical bound allows, at the necessary expense of the mixed outcomes (++- and +-+).
This deviation is a direct signature of contextuality. Unlike the classical model, the quantum system is not constrained to maintain consistency with unperformed measurements. The violation of Bell-type inequalities thus serves a dual purpose: it certifies the privacy of the time stream against a middleman adversary, and it confirms that the sovereign quantum clock ticks away faster than allowed by the constraints of classical reality.